OpenStudy (anonymous):

A recipe for oatmeal raisin cookies call for 1 2/3 cups of flour to make 4 dozen cookies. How many cups of flour are needed to make 6 dozen cookis?

OpenStudy (anonymous):

well, you can see that if 5/3 cups make 4, that means you need to add half of that to make 2 more dozen (6-4 = 2) so, (5/3)(1/2) = 5/6 which you add to the original 5/3 5/3 + 5/6 = 10/6 + 5/6 = 15/6 cups or 2 1/2

OpenStudy (anonymous):

Notice that going from 4 dozen to 6 dozen is increasing the recipe by 50%. So the original recipe calls for 1-2/3 cups of flour. You can convert this to 5/3 cups of flour. So you want to increase 5/3 by 50%. That would be: (5/3) X (3/2) That gives you 5/2 cups; that's the same as 2-1/2 cups. Make sense? Good luck!

OpenStudy (anonymous):

thanks guys:D

OpenStudy (anonymous):

glad to help =)